OpenStudy (anonymous):

solve for A: A^2+4=45

OpenStudy (anonymous):

answer is square root of 41

OpenStudy (anonymous):

\[a = - \sqrt{41}\] \[a = + \sqrt{41}\]

OpenStudy (jwt625):

sqt41

OpenStudy (anonymous):

\[A^2+4-4=45-4\] \[A^2=41\] \[A=\pm \sqrt{41}\]